OpenStudy (anonymous):

hello, can someone please explain to me how the integral of sec^2(x) / tan(x) dx, = ln tan(x) + c ? I don't understand how to simplify sec^2(x)/tan(x) to something that integrates into that. thanks very much

OpenStudy (lalaly):

u= tanx du=sec^2 x dx so integral of du/u= lnu since u is tanx so ln tanx

OpenStudy (lalaly):

ln tanx +c ofcourse...i forgot that

OpenStudy (anonymous):

Recall that: \[\cos^2 x + sin^2 x = 1\]Dividing by \(cos^2 x\)\[ \implies1 + tan^2x = sec^2 x\] \[\implies \int {sec^2x \over tan x} dx = \int {1+tan^2x \over tan x}dx\]

OpenStudy (anonymous):

\[ = \int [\frac{1}{tanx} + tanx] dx\] Which you can integrate term by term.

OpenStudy (anonymous):

Or you can use a u-sub. That way is probably faster

OpenStudy (anonymous):

thank you both.